# Representational Stability of Truth in Large Language Models

### v2

@@ 4.1 Data
Our experiments draw on the three factual domains introduced in [savcisens2025trilemma]: City Locations, Medical Indications, and Word Definitions. Although all three domains contain factual assertions, they differ in how sharply truth and falsehood are delineated. Statements in City Locations are objective and stable. Statements in Medical Indications are factual but context-dependent. Statements in Word Definitions are more interpretive due to polysemy and variation in usage. This range provides a diverse testbed for studying how LLMs encode veracity across domains. Full dataset construction and validation details appear in Supplementary Section B.
We analyze five types of statements: True, False, Synthetic, Fictional, and Noise (Table 1). We take the True, False, and Synthetic statements directly from [savcisens2025trilemma]. Synthetic statements are grammatically coherent but semantically meaningless constructions built from generated entity names. Because these entities cannot have appeared in training corpora, LLMs lack the background needed to assign them a truth value. They therefore serve as unfamiliar Neither statements for which calibrated models should suspend belief.
Fictional statements lack real-world truth value. [^1] We have released the fictional statements at https://huggingface.co/datasets/samanthadies/representational_stability. Unlike unfamiliar Synthetic statements, many fictional entities (e.g., Gotham City, Xenovirus Takis-B) are likely present in training corpora. Fictional statements test whether LLMs distinguish between recognition and factual commitment: a model may have rich associations with these entities while still treating them as nonfactual.
Finally, Noise serves as a non-semantic control. We sample random activation sequences matched to the mean, variance, and sequence-length distribution of the real activations. This ensures that observed representational effects arise from semantic content and context rather than from activation-space statistics alone.
4.2 Stability
We measure stability to discuss the probe, the activations, and the perturbations.
Probe.
We use the sparse-aware multiple-instance learning probe (sAwMIL) [savcisens2025trilemma], a multiclass probing method designed to extract reliable and transferable veracity directions from LLM activations. Unlike simpler probes such as the Mean Difference classifier [marks2310geometry], which assumes that truth and falsehood lie along a single axis, sAwMIL models True, False, and Neither as distinct directions and aggregates token-level representations using multiple-instance learning. It also incorporates conformal prediction [angelopoulos2023conformal] to calibrate uncertainty. As a max-margin method, sAwMIL yields stable decision boundaries, making differences across perturbations more reflective of genuine structure in the LLM’s geometry than probe noise. For comparison, results from the Mean Difference probe appear in Supplementary Section E.
Generating and Characterizing Activations.
@@ 4.2 Stability
Figure 3: Average Wasserstein distance between activations. Pairwise Wasserstein distances between activation distributions of True, False, Synthetic, Fictional, and Noise statements, averaged over sixteen LLMs for the (a) City Locations, (b) Medical Indications, and (c) Word Definitions datasets. Across datasets, Synthetic activations lie closest to the True and False activations, while Fictional and Noise activations are farther from all others, indicating that unseen but fact-like statements (Synthetic) resemble factual structure, whereas Fictional statements form distinct representational clusters.
@@
For descriptive, model-agnostic analyses, we reduce each sequence to a single vector by selecting the final non-padding token. We then characterize patterns at both the linguistic and representation levels. At the linguistic level, we compute rank–frequency curves over character bigrams aggregated across entity names for each statement type. At the representation level, we compute pairwise $1$-D Wasserstein distances between activation distributions across all dimensions, averaged across LLMs. These metrics reveal similarities between statement types at the linguistic level and in latent space before any supervised probing.
Perturbations.
[…]
Experimental Details.
@@
Activation sequences, data splits, hyperparameters, and preprocessing steps are held fixed. Token-level representations are scaled using a standard scaler fit on the training set; bags are truncated to a fixed maximum size; and we perform a small grid search over the regularization parameter $\mathcal{C}$ using three-fold cross-validation with mean average precision as the criterion. Approximately $55\%$ of the data is used for training, $20\%$ for calibration, and $25\%$ for testing (see Supplementary Table A2). Holding all training conditions constant ensures that changes in the learned truth direction arise solely from label perturbations.
Figure 4: Changes in the probe decision boundary under perturbations. Cosine similarity (left column) and bias difference (right column) between the baseline True vs. Not True probe and probes retrained under label perturbations for the (a,b) City Locations, (c,d) Medical Indications, and (e,f) Word Definitions datasets. Each heatmap shows results for sixteen LLMs (columns) and five perturbation conditions (rows). LLMs with leading underscores are chat models, while those without are base models. Higher cosine similarity indicates smaller rotations of the learned decision boundary, while bias difference reflects shifts in intercept. Across datasets, probes retrained with the Synthetic perturbation show the largest deviation from the original, particularly in cosine similarity.
Table 3: Flipped predictions under label perturbations. Counts (and percentages) of predictions that remain stable or flip between True and Not True across Synthetic, Fictional, Fictional(T), and Noise perturbations for each dataset. City Locations is the most stable, with only $4.8\%$ of statements flipping in the worst case, followed by Medical Indications ($12.2\%$) and Word Definitions ($40.6\%$). Across all domains, Synthetic perturbations produce the most flips, suggesting that unseen but fact-like statements most strongly distort the learned veracity boundaries.
Dataset Perturbation True to True Not True to Not True Not True to True True to Not True Total Flips
City Locations Synthetic $9153$ $(91.5)$ $360$ $(3.6)$ $274$ $(2.7)$ $213$ $(2.1)$ $\mathbf{460}$ $\mathbf{(4.8)}$
Fictional $9326$ $(93.3)$ $568$ $(5.7)$ $66$ $(0.7)$ $40$ $(0.4)$ $106$ $(1.1)$
Fictional (T) $9330$ $(93.3)$ $576$ $(5.8)$ $58$ $(0.6)$ $36$ $(0.4)$ $94$ $(1.0)$
Noise $9183$ $(91.8)$ $532$ $(5.3)$ $102$ $(1.0)$ $183$ $(1.8)$ $285$ $(2.8)$
Medical Locations Synthetic $7413$ $(72.6)$ $1556$ $(15.2)$ $786$ $(7.7)$ $460$ $(4.5)$ $\mathbf{1246}$ $\mathbf{(12.2)}$
Fictional $7808$ $(76.4)$ $2284$ $(22.4)$ $58$ $(0.6)$ $65$ $(0.6)$ $131$ $(1.3)$
Fictional (T) $7815$ $(76.5)$ $2269$ $(22.2)$ $73$ $(0.7)$ $58$ $(0.6)$ $427$ $(1.3)$
Noise $7779$ $(76.2)$ $2009$ $(19.7)$ $333$ $(3.3)$ $94$ $(0.9)$ $427$ $(4.2)$
Word Definitions Synthetic $3682$ $(37.2)$ $2188$ $(22.1)$ $785$ $(7.9)$ $3233$ $(32.7)$ $\mathbf{4018}$ $\mathbf{(40.6)}$
Fictional $6507$ $(65.8)$ $2494$ $(25.2)$ $479$ $(4.1)$ $408$ $(4.1)$ $887$ $(8.2)$
Fictional (T) $6795$ $(68.7)$ $2520$ $(25.5)$ $453$ $(4.6)$ $120$ $(1.2)$ $573$ $(5.8)$
Noise $6653$ $(67.3)$ $2251$ $(25.4)$ $462$ $(4.7)$ $262$ $(2.6)$ $724$ $(7.3)$
Figure 5: Stability of probe predictions under label perturbations for City Locations data. Bar plots show, for each of the sixteen LLMs (x-axis), how often the sAwMIL probe’s predicted label changes when retrained under four perturbations: (a) Synthetic, (b) Fictional, (c) Fictional(T), and (d) Noise. Green bars indicate True to Not True flips, while purple bars indicate Not True to True flips. The left y-axis reports the number of statements with flipped predictions, and the right y-axis reports the corresponding proportions. The Synthetic perturbation leads to the most instability, and the base models exhibit more True to Not True flips than the chat models.
Figure 6: Stability of probe predictions under perturbations for Medical Indications data. Bar plots show, for each of the sixteen LLMs (x-axis), how often the sAwMIL probe’s predicted label changes when retrained under four perturbations: (a) Synthetic, (b) Fictional, (c) Fictional(T), and (d) Noise. Green bars indicate True to Not True flips, while purple bars indicate Not True to True flips. The left y-axis reports the number of statements with flipped predictions, and the right y-axis reports the corresponding proportions. The Synthetic perturbation leads to the most instability, and the Fictional and Fictional(T) perturbations result in almost no flips.
Figure 7: Stability of probe predictions under label perturbations for Word Definitions data. Bar plots show, for each of the sixteen LLMs (x-axis), how often the sAwMIL probe’s predicted label changes when retrained under four perturbations: (a) Synthetic, (b) Fictional, (c) Fictional(T), and (d) Noise. Green bars indicate True to Not True flips, while purple bars indicate Not True to True flips. The left y-axis reports the number of statements with flipped predictions, and the right y-axis reports the corresponding proportions. The Synthetic perturbation leads to the most instability, with some LLMs retracting over $50\%$ of their originally True statements.
@@ 5 Results
We present results at three levels of analysis. First, we examine the input representations: how statement types differ in linguistic-level features and in activation space. Second, we measure probe-level changes by assessing how the learned True vs. Not True direction shifts under controlled label perturbations. Third, we analyze output-level changes in the probe’s predicted labels, quantifying how truth-value assignments respond to modified semantic assumptions.
5.1 Representations of True, False, and Neither Statements
@@
We first characterize the probe inputs. At the linguistic level, True, False, and Synthetic statements exhibit nearly identical normalized bigram distributions across all three domains (Figure 2). This confirms that the generation procedure for Synthetic statements [^3] See Supplementary B.1.1 for details on the procedure that generates Synthetic statements. preserves low-level linguistic structure. Fictional statements, however, show a slower rank–frequency decay, reflecting stylistic patterns characteristic of narrative text. This divergence is most visible in Word Definitions (Fig. 2(c)) and least visible in City Locations (Fig. 2(a)), where some fictional cities align with real countries (e.g., “Brigadoon is located in Scotland”).
We next test whether these linguistic-level differences correspond to differences in latent space by computing pairwise Wasserstein distances between activation distributions for all statement types (Fig. 3; per-model heatmaps appear in Supplementary Figures A1A16).
Across all sixteen LLMs, True and False representations lie close together, with mean distances of $0.40$, $0.29$, $0.14$ for City Locations, Medical Indications, and Word Definitions, respectively.
Synthetic statements remain only modestly farther from True, with mean distances of $0.58$, $0.44$, $0.33$ for City Locations, Medical Indications, and Word Definitions, respectively. This is consistent with their linguistic similarity despite their unfamiliarity.
By contrast, Fictional and Noise statements lie much farther from True. For Fictional, the mean distances from True are $2.28$, $1.30$, $1.41$ for City Locations, Medical Indications, and Word Definitions, respectively. For Noise, the mean distances from True are $1.37$, $1.30$, $1.09$ for City Locations, Medical Indications, and Word Definitions, respectively. The mean distances between Fictional and Noise are $2.38$, $1.93$, $1.58$ for City Locations, Medical Indications, and Word Definitions, respectively. Fictional statements form their own representational cluster shaped not only by lexical differences but by their appearance in non-factual contexts during training.
@@
Taken together, these findings show a clear decoupling between linguistic and latent-space similarity. Word Definitions exhibit substantial linguistic divergence between factual and fictional content, yet only moderate activation-space divergence. Conversely, City Locations show minimal linguistic differences but large representational separation. The geometry of LLM activations thus reflects both linguistic form and epistemic context.
5.2 Probe-level Changes under Label Perturbations
We assess representational stability by retraining the probe with expanded definitions of True (i.e., by treating Synthetic, Fictional, Fictional(T), or Noise as True). Because the LLM activations remain fixed, shifts in the learned boundary reflects how the underlying veracity structure supports (or resists) linear reclassification.
Figure 4 reports cosine similarity and intercept shifts between baseline and perturbed classifiers. Cosine similarity captures rotations of the True vs. Not True direction, while bias differences capture translational shifts. Across the datasets, Synthetic perturbations induce the largest boundary rotations, often nearing orthogonality in Word Definitions (Fig. 4(e),(f)). Fictional and Noise perturbations yield considerably smaller deviations. These results indicate that the baseline True vs. Not True direction is generally stable, but that adding unfamiliar yet semantically factual statements forces substantial reorientation. Synthetic statements therefore reveal where the veracity structure is most brittle.
Although the global pattern is consistent, different LLM families exhibit different degrees of susceptibility to these perturbations. Chat-tuned variants (denoted by leading underscores) tend to exhibit somewhat larger rotations and bias shifts than their base models. An exception is gemma-7b, which shows unusually large shifts in the Word Definitions domain. Importantly, such rotations do not necessarily imply failures of predictive accuracy: large re-orientations can reflect either weak separation or greater dispersion within the underlying latent geometry (e.g., models with larger inter-type Wasserstein distances). Overall, across models and domains, Synthetic perturbations consistently produce the strongest probe-level instability.
5.3 Changes in Predicted Labels
We examine how shifts in the truth boundary affect predicted veracity labels. Table 3 reports the percentage of statements whose predicted labels remain stable versus those that flip between True and Not True. City Locations is most stable (maximum flip rate $4.8\%$), Medical Indications shows intermediate instability ($12.2\%$), and Word Definitions is substantially less stable (up to $40.6\%$). In all domains, Synthetic perturbations yield the highest flip rates, reinforcing that unfamiliar yet semantically fact-like statements are most disruptive to the learned veracity boundary.
Figures 57 show flip patterns by model family. Chat-tuned models tend to produce more Not True to True flips, suggesting a mild tendency toward over-inclusiveness when the True class expands. Base models tend to show the opposite pattern. From a $P$-stability perspective, True to Not True flips represent stronger epistemic instability, as they retract previously assigned beliefs. However, these tendencies do not hold uniformly, and overall differences across LLM families are smaller than differences across perturbation types.
These output-level results reveal a consistent hierarchy of representational stability. Domains richly represented in training corpora (e.g., geography) exhibit the strongest representational stability; specialized but widely discussed domains (e.g., medicine) show moderate stability; and semantically flexible domains (e.g., definitions) are most vulnerable. Across all cases, Synthetic perturbations remain the dominant source of instability, indicating that unfamiliar yet factually plausible statements impose the greatest challenge for generating well-organized truth representations in LLMs.
@@ 6 Discussion
Our results show that LLMs encode a coherent but sometimes brittle separation of veracity in their internal activations. True, False, and familiar Fictional statements form well-defined clusters, while unfamiliar Synthetic statements sit near the boundary and disrupt it when relabeled as True. Because the activations remain fixed, these disruptions reflect genuine weaknesses in the underlying veracity geometry rather than artifacts of training or optimization. The mismatch between linguistic similarity and activation-space similarity across domains further indicates that representational stability is driven by epistemic familiarity, not linguistic-level features.
A central implication is that representational stability depends strongly on training-induced familiarity. Familiar Fictional statements appear in narrative contexts the LLMs have likely repeatedly encountered, enabling them to form stable, internally coherent clusters. Unfamiliar Synthetic statements lack any such anchoring: they resemble factual claims but violate the LLM’s learned priors. As a result, they induce the largest rotations and label flips, mirroring how humans show greater belief instability when confronted with novel but superficially plausible assertions. In this way, the probe’s decision boundary serves as a diagnostic of the model’s epistemic landscape, revealing which statements the model encodes as grounded, merely recognized, or unsupported.
Domain differences reinforce this picture. City Locations, which are richly represented and highly regular in training data, shows the tightest clustering and highest stability. Medical Indications, which are factual but context-sensitive, shows moderate stability. Word Definitions, which are semantically flexible and often usage-dependent, show the weakest structure. This reflects broader patterns in factual generalization: LLMs inherit the uneven epistemic structure of their training corpora, and their capacity to encode veracity varies accordingly.
Beyond characterizing these patterns, our method helps diagnose the representational sources of factual inconsistency. By perturbing the labeling of truth rather than modifying LLM parameters, we isolate which semantic shifts the underlying geometry will tolerate. This complements output-based factuality metrics: rather than asking whether a model states the truth, we evaluate whether it represents truth in a stable, well-organized fashion. Such diagnostics could guide data curation, fine-tuning objectives, or auditing procedures focused on improving epistemic reliability rather than linguistic-level accuracy.
Finally, our results connect to philosophical accounts of $P$-stability [leitgeb2014stability], which hold that rational beliefs should remain stable under small evidential changes. In our setting, reclassifying Neither statements that are logically consistent with the True statements should preserve the inferred truth boundary if the model’s epistemic representation is robust. Instead, Synthetic perturbations cause substantial reorientation. This highlights a deeper challenge: distributional similarity alone does not guarantee that truth, falsity, and indeterminacy are encoded in forms that support stable inference. Addressing this may require training objectives or architectures that explicitly distinguish these epistemic categories.
Future Work.
Our analysis focuses on representational (and not behavioral) stability. We evaluate how truth directions emerge from fixed activations rather than how LLMs revise beliefs. Extending the framework to dynamic settings, such as after fine-tuning, reinforcement learning, or conversational interaction, may reveal how representational shifts propagate to behavior. Although sAwMIL [savcisens2025trilemma] provides a strong linear diagnostic, nonlinear or causal probes could uncover subtler dependencies among veracity, uncertainty, and linguistic form. Finally, our datasets center on factual and quasi-factual content; applying the method to disputed, opinion-based, counterfactual, or evolving claims would broaden the range of epistemic ambiguity and deepen our understanding of how LLMs encode belief.

### v3

@@ 4 Experiments
We use P-StaT to apply the same perturbation $\Theta$ to both (i) probe-based evaluations over $\phi_{\mathcal{M},l}$ and (ii) zero-shot evaluations via belief context. We implement all experiments in Python using PyTorch [21], NumPy, SciPy, and scikit-learn [22], and rely on the HuggingFace Transformers ecosystem for LLM access and data handling [34]. All experiments were run on a university HPC cluster with NVIDIA H200 GPUs and required approximately $36$ GPU-hours (activation extraction and zero-shot evaluations), with probe training performed on CPU.
@@ 4.1 Data
We use the three domains and statement types defined in Section 3.1 (Table 1). Fictional is newly released with this work, and Noise is a set of randomly generated Gaussian vectors that match the dimensionality and distribution of the veracity representations and serve as a probing-only non-semantic control. Supplementary Section B contains additional information on data construction.
Approximately $55\%$ of the data is used for training, $20\%$ for calibration, and $25\%$ for testing (see Supplementary Tab. A2). The splits are shared across all experiments. We compute stability on the same held-out set of ground-truth True statements from the test split, so that differences in retractions are attributable only to changes in $\Theta$.
4.2 Activations
@@
For descriptive analyses, we reduce each activation sequence to a single vector by selecting the final non-padding token. At the linguistic level, we compute rank–frequency curves over character bigrams aggregated across entity names for each statement type. At the representation level, we compute pairwise $1$-D Wasserstein distances between activation distributions, considering all activation dimensions. These measures reveal similarities between statement types at the linguistic level and in latent space.
4.3 Perturbation Conditions and Shared Protocol
For each perturbation $\Theta$ in Table 2, we instantiate $\Theta$ using only training data: probes are retrained with labels induced by $\Theta$, and belief contexts $C_{\Theta}$ are constructed from the corresponding training statements. We then evaluate on the same held-out set of ground-truth True test statements, computing $\mathcal{B}^{\Theta_{0}}_{\mathrm{true}}$ and $\mathcal{B}^{\Theta}_{\mathrm{true}}$ by applying $g(\cdot,\Theta_{0})$ and $g(\cdot,\Theta)$, and report epistemic retractions $\mathcal{R}=\mathcal{B}^{\Theta_{0}}_{\mathrm{true}}\setminus\mathcal{B}^{\Theta}_{\mathrm{true}}$ (and expansions analogously).
4.3.1 Instantiation I: Probing over Activations
To implement $g(\cdot,\Theta)$ representationally, we train a linear probe that operates on the veracity representations $\phi_{\mathcal{M},l}(s)$. We use the sparse-aware multiple-instance learning probe (sAwMIL) [23], a multiclass probing method designed to extract reliable and transferable veracity directions from LLM activations. Unlike simpler probes such as the Mean Difference classifier [20], which assumes that truth and falsehood lie along a single axis, sAwMIL models True, False, and Neither as distinct directions and aggregates token-level representations using multiple-instance learning. As a max-margin method, sAwMIL yields stable decision boundaries, making differences across perturbations more reflective of genuine structure in the LLM’s geometry than noise from the probe. [^2] For comparison, results from the Mean Difference probe appear in Supplementary Section E.
@@ 4.3.1 Instantiation I: Probing over Activations
For each condition in Table 2, we retrain the probe on $\mathcal{D}_{\mathrm{train}}$ with labels induced by $\Theta$, holding the architecture and hyperparameters fixed across conditions. Token-level representations are scaled using a standard scaler fit on the training set; bags are truncated to a fixed maximum size; and we perform a grid search over the regularization parameter $\mathcal{C}$ using three-fold cross-validation with mean average precision. [^3] Our code with all seeds and final hyperparameters is at https://github.com/samanthadies/representational_stability.
4.3.2 Instantiation II: Zero-shot Prompting with Belief Context
To implement $g(\cdot,\Theta)$ behaviorally, we use a zero-shot prompt. For each condition $\Theta$ (Tab. 2), we construct a belief context $C_{\Theta}$ from the corresponding training statements and insert it into each held-out ground-truth True test statement $s$.
To construct $C_{\Theta}$, we uniformly sample $K=100$ context statements from $\mathcal{N}_{\Theta}$ as specified in Table 2. We sample without replacement.
For each ground-truth True test statement, we use the following prompt:
[optional belief context $C_{\Theta}$]
Is the following statement correct?
 [statement $s$]
a. The statement is true.
b. The statement is false.
c. The statement is neither true nor false.
The final answer is
For chat-tuned LLMs, we place the same content into the LLM’s chat template and define the model’s response as the highest-probability next token among the answer labels $\{a,b,c\}$ at the final prompt position.
We compute next-token probabilities for $a$, $b$, and $c$ at the final position and predict
$g_{\mathrm{zs}}(s,\Theta)=\mathop{\mathrm{arg\,max}}\limits_{\ell\in\{a,b,c\}}p(\ell\mid s,\Theta),$
with temperature set to zero. [^4] We also track the probability mass corresponding to all non-label tokens, but define $g$ exclusively over $\{a,b,c\}$ rather than parsing free-form textual outputs. We treat option $a$ as True and options $\{b,c\}$ as Not True. All zero-shot evaluations are therefore deterministic, and reported results aggregate outcomes across LLMs rather than multiple random seeds.
@@ 5 Results
Figure 2: Linguistic vs. representational structure of Neither statements. For (a),(b) City Locations, (c),(d) Medical Indications, and (e),(f) Word Definitions, the left column shows normalized character bigram rank–frequency curves for True (green), False (red), Synthetic (yellow), and Fictional (blue) statements, and the right column shows pairwise Wasserstein distances between activation distributions. Fictional content exhibits a domain-dependent decoupling between linguistic form and latent-space organization, indicating that veracity representations reflect both epistemic context and linguistic form.
Table 3: Epistemic expansions $\mathcal{E}$ and retractions $\mathcal{R}$ under probing label perturbations. Counts (percentages) of beliefs that remain stable or undergo expansions $\mathcal{E}$ or retractions $\mathcal{R}$ under Synthetic (yellow), Fictional (gray), Fictional(T) (blue), and Noise (red) perturbations. Synthetic perturbations consistently produce the highest rate of epistemic retractions, indicating that epistemically unfamiliar content induces the most instability.
Table 4: Epistemic expansions $\mathcal{E}$ and retractions $\mathcal{R}$ under perturbed zero-shot experiments. Counts (percentages) of beliefs that remain stable or undergo expansions $\mathcal{E}$ or retractions $\mathcal{R}$ under Synthetic (yellow), Fictional (gray), and Fictional(T) (blue) perturbations. Despite domain variation, Synthetic perturbations induce the highest retraction rates in all cases, mirroring the hierarchy observed under probing in Table 3.
We report (i) the descriptive structure of Neither statements, (ii) representational stability under probing, and (iii) behavioral stability under zero-shot prompting. LLM-level results are in Supplementary Section D.
5.1 Linguistic and Representational Structure of Neither Statements
@@
We begin by characterizing how familiar and unfamiliar Neither statements differ. At the linguistic level, True, False, and Synthetic statements exhibit nearly identical normalized bigram rank–frequency curves (Figure 2(a),(c),(e)). Fictional statements, by contrast, show a slower decay reflecting stylistic patterns characteristic of narrative text. The differences in Fictional bigram distributions are most visible in Word Definitions and least in City Locations.
We next test whether these linguistic differences explain latent-space structure by comparing activation distributions across statement types using pairwise Wasserstein distances (Figure 2(b),(d),(f); LLM-level heatmaps appear in Supplementary Section D.1). Across domains, True and False occupy nearby regions of activation space, and Synthetic remains relatively close to factual content despite its lack of real-world referents. By contrast, Fictional statements are consistently more representationally separated from factual content, forming a distinct cluster that is not well-explained by bigram statistics alone.
@@
Taken together, these findings indicate a decoupling between linguistic and latent-space similarity. Although linguistic divergence of Fictional content varies by domain, representational separation remains consistently pronounced. Further, the domain with the largest Fictional separation linguistically (Word Definitions) is the domain with the smallest representational distance. The geometry of LLM activations thus reflects both linguistic form and context.
5.2 Representational Stability under Probing Perturbations
We next evaluate P-StaT representational stability by retraining sAwMIL under perturbed interpretations $\Theta$ and measuring epistemic retractions on held-out ground-truth True statements. Table 3 summarizes prediction changes across perturbation conditions aggregated over all LLMs. LLM-level results and results for the Mean Difference probe appear in Supplementary Sections D.2 and E.
Across all three domains, Synthetic perturbations induce the highest rate of epistemic retractions: $2.1\%$ in City Locations, $4.5\%$ in Medical Indications, and $32.7\%$ in Word Definitions. These rates are substantially larger than those produced by Fictional and Fictional(T) perturbations, establishing a clear perturbation hierarchy in which epistemically unfamiliar Neither content most strongly destabilizes learned veracity boundaries. Synthetic perturbations also induce the largest epistemic expansions across domains.
Retraction rates also vary systematically by domain. City Locations is highly stable, Medical Indications shows moderate instability, and Word Definitions is markedly more fragile. These results again indicate that representational stability is governed by epistemic familiarity, given that City Locations are likely richly represented in training corpora, while Medical Indications are common but specialized, and Word definitions are the most semantically flexible.
5.3 Behavioral Stability under Zero-Shot Perturbations
We next evaluate P-StaT behavioral stability by applying the same perturbations $\Theta$ via belief context in zero-shot prompting and measuring epistemic retractions. Table 4 reports belief changes across LLMs, with LLM-level results in Supplementary Section D.3. Overall belief changes are larger than in probing experiments, reflecting the greater flexibility of behavioral responses compared to linear reclassification over fixed representations.
Despite this increase, the perturbation hierarchy observed representationally also appears behaviorally. The Synthetic perturbation induces the highest epistemic retraction rate in every domain, with $36.3\%$ in City Locations, $27.1\%$ in Medical Indications, and $21.4\%$ in Word Definitions. Fictional and Fictional(T) also induce substantial retractions, but consistently fewer than Synthetic within each dataset.
However, domain ordering differs from the probing setting. City Locations has the highest zero-shot retraction rates, while Word Definitions has the lowest. This reversal likely reflects interactions between belief stability and behavioral accuracy. Specifically, zero-shot retractions conflate belief changes with the number of baseline beliefs available to retract (i.e., a statement must be considered True in the baseline case in order to become a retraction). In contrast, probing isolates changes in the decision boundary under fixed representations.
Our results show that epistemic familiarity governs stability not only in activation space, but also in LLM behavior. Unfamiliar Synthetic content systematically destabilizes LLM beliefs under semantic perturbation, whereas familiar Fictional content produces smaller and more context-dependent effects.
@@ 6 Discussion
We provide a unified perspective on how LLMs organize and maintain truth judgments under semantic perturbation. Across LLMs, domains, and evaluation settings, Neither statements play a central role in belief stability: unfamiliar Synthetic statements consistently induce the largest epistemic retractions, while familiar Fictional statements are more stable. Our findings indicate that belief stability in LLMs is governed as much by epistemic familiarity as by linguistic form.
At the representational level, we observe a decoupling between linguistic similarity and latent-space organization. Fictional content exhibits domain-dependent linguistic variation that does not align with distance in representation space. This mismatch suggests that veracity representations encode higher-level epistemic context learned during training rather than lexical statistics alone. Through the lens of P-StaT, these representational differences translate into systematic differences in stability under semantic perturbation. Expanding the semantic definition of truth to include unfamiliar Synthetic content reliably destabilizes previously held beliefs, while perturbations involving familiar Fictional content are better tolerated. These results hold in both representational and behavioral instantiations, suggesting that latent-space organization captures meaningful epistemic structure.
More broadly, P-StaT complements accuracy-based factuality benchmarks by shifting focus from correctness under fixed prompts to stability under principled semantic shifts. By emphasizing epistemic retractions, it targets a particularly consequential form of instability. This perspective aligns with formal accounts of rational belief change, such as Leitgeb’s notion of $P$-stability, which requires that established beliefs be preserved under justified changes in evidential context [17]. From this viewpoint, the sensitivity of LLMs to Synthetic perturbations highlights a structural limitation: distributional plausibility alone does not ensure a stable internal organization of truth, falsity, and indeterminacy. Together, these findings suggest that evaluating belief stability under controlled semantic perturbations offers a principled way to probe the epistemic organization of LLMs beyond what can be inferred from accuracy alone.